Summary of Second Lecture on Dasein

Greg Hudson

November 18, 1994

In this lecture, we define some new terms in order to study the phenomenon of authenticity. The first term is “the already”, the existing state of a project, as opposed to the beyond, the goals of the project. The second term is “das man”, the impersonal other, which brings us customs and comforts as well as the stifling oftradition. The third term is “fallenness,” or refusing to listen to the call of the beyond; this is similar to the natural attitude we looked at before. The fourth term is “authenticity,” or heeding the call of the beyond.

Some of the things we must do as phenomenologists: we must consider psychological phenomena as objective phenomena, rather than something “in” the mind; we must describe learning using the Hermeneutic circle (repetition leading to familiarity) rather than as traditional deductive logic, and we must consider truth as an ontic instance of evidence rather than as an absolute rubber stamp (since a rubber stamp would lead to an infinite regression, as we try to certify the rubber stamp).

An example of fallenness and authenticity is when Professor Rota tries to solve a math problem. Rather than face the problem authentically, he gathers literature on the problem and studies it, comes up with solutions which are obviously wrong but which he might convince himself are correct. Finally, he authentically faces the problem and sees the solution, and suddenly feels that the inauthentic steps were unnecessary. At the same time, there is a second fall, as the solution to the problem falls back into the everyday dealings with the world.

“Mitsein” is the term we use for “co-being” or “being with.” In reality, all being is co-being, and we must always be prepared to consider being as a shared phenomenon.